OpenStudy (anonymous):

what is 5 to the 2nd power times 2to the 3td power

OpenStudy (ageta):

5^2 *2^3

OpenStudy (anonymous):

\[5^{2} \times 2 ^{3} = \]

OpenStudy (anonymous):

yes but what is the answer to that question

OpenStudy (anonymous):

\[Multiply~5 \times 5. ~Then~Multiply~ 2\times2\times2~To~Get~Your~Anwser. \]

OpenStudy (anonymous):

\[5 \times 5 = 25 \]

OpenStudy (anonymous):

\[2 \times 2 \times 2 = 8 \]

OpenStudy (skullpatrol):

25 X 8 = ?

OpenStudy (skullpatrol):

(20 + 5)8 = 20 X 8 + 5 X 8 = 160 + 40 =?

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

Multiply 25 x 8 to get your anwser.

OpenStudy (anonymous):

Do you know what 25 x 8 is ?

OpenStudy (skullpatrol):

The distributive property is your friend :)

OpenStudy (anonymous):

no it is not!

OpenStudy (skullpatrol):

:O

OpenStudy (anonymous):

Do you still need help or are you okay with this question now?

OpenStudy (anonymous):

I am ok thanks!

OpenStudy (skullpatrol):

$$\Huge\text{a(b+c)=ab+ac}$$

OpenStudy (anonymous):

Alright. You're Welcome.

OpenStudy (anonymous):

I gave you a medal for helping

OpenStudy (anonymous):

Thanks!

OpenStudy (skullpatrol):

:(

OpenStudy (anonymous):

is rachella your real name

OpenStudy (anonymous):

I gave you a medal @skullpatrol for your willingness to help!

OpenStudy (anonymous):

No, my name is Rachel. Rachella is just a nickname.

OpenStudy (anonymous):

my name is jazmin and thanks for the medal and how do you become a fan

OpenStudy (anonymous):

wait I think I know it now!

OpenStudy (anonymous):

Nice to meet you Jazmin! To become a fan, you simply click on the user that you want to fan and click 'become a fan.'